OpenStudy (konradzuse):

Find the vertex, focus, and directrix of the parabola y^2+6y+4x+1=0

OpenStudy (konradzuse):

so far I have y^2 + 6y = -4x-1 = 4(-x+1)

OpenStudy (konradzuse):

but cannot factor y^2+6y.... SO I might need to go by a diff approach.

OpenStudy (anonymous):

okay complete the square for the left side. by adding and subtracting 9 on left.

OpenStudy (konradzuse):

Ha funny because I just entered that into wolfram to see if it would factor :P. (y^2 + 6y +9) -9 = .... Y+3)^2 - 9 = ....

OpenStudy (konradzuse):

sorry openstudy just crashed me :(.

OpenStudy (anonymous):

okay now you have (y+3)^2=-4x+8 (+3)^2=-4(x-2) compare with standard form (y-b)^2=4c(x-a) where (a,b) is vertex and c is the distance between focus and vertex. hope you can do this now.

OpenStudy (konradzuse):

Thanks, I can finish this now :D

OpenStudy (anonymous):

you 're welcome :)

OpenStudy (konradzuse):

Very useful that formula... Can you tell me the ones for ellipse, and hyperbola too please?

OpenStudy (konradzuse):

Oh and how do we find x for the focus? c?

OpenStudy (konradzuse):

in this case x = 1, but in another case c = 1/4 and the x was 0.

OpenStudy (anonymous):

yeah sure. for ellipse \[\Large \frac{(x-a)^2}{h^2}+\frac{(y-b)^2}{k^2}=1\] it is centered at (a,b) and h,k are length of major and minor axis. for hyperbola \[\Large \frac{(x-a)^2}{h^2}-\frac{(y-b)^2}{k^2}=1\] (a,b ) is center and h,k are major and minor axis respectively.

OpenStudy (anonymous):

x for the focus is found by first find distance c -4c=4 c=-1 |c|=1 since vertex is (2,-3) subtract c from the x coordinate (1,-3) is focus.

OpenStudy (konradzuse):

hmmm.. ok gotta read over this thanks :).

OpenStudy (konradzuse):

oic that makes sense thanks :D.

OpenStudy (konradzuse):

and how do we find the directrix?

OpenStudy (konradzuse):

I got it as x = 3, it seems like it's just the value of y.

OpenStudy (konradzuse):

so 2--1 = 3? :)

OpenStudy (anonymous):

directrix is just at the same distance from vertex as focus is but is at the right side of vertex so since C=1 so x=1 is directrix

OpenStudy (konradzuse):

from the example and what you said it seems like if your'e doing -c for the focus and c for directrix... In the example it's -1/4 so if you do +1/4 you get 0 which is correct for the focus... But if you do -1/4 - 1/4 you get -2/4 which is -1/2 directrix which is what the correct answer is SO Idk..

OpenStudy (anonymous):

sorry the directx is x=3 because 1 should be added to x coordinate . x=3

OpenStudy (konradzuse):

so it's x + c?

OpenStudy (anonymous):

yes but remember c is distance so take it as positive..

OpenStudy (konradzuse):

hmm?

OpenStudy (konradzuse):

So how does it work in the other case?

OpenStudy (anonymous):

x+c is equation of directx . c is positive.

OpenStudy (konradzuse):

y2 − x − y = 0

OpenStudy (konradzuse):

Vertex is (-1/4,1/2) Directix is x = -1/2

OpenStudy (konradzuse):

c = 1/4 if it's -1/4 + 1/4 it should be 0... which is the focus x.

OpenStudy (konradzuse):

Maybe I'm just confused haha.

OpenStudy (anonymous):

you should be able to do this now !!! you should look at the graph .if it opens leftward then add c to x coordinate. if opens rightward subtract c from x coordinate. just look at the graph dirextx should always be behind the vertex at the distance c . by looking to graph make appropriate decision.

OpenStudy (konradzuse):

oic :).

OpenStudy (konradzuse):

One more thing how do we find the vertices of a parabola?